Part II: Solve the following problem

1.      Consider in perfectly competitive market the following demand and supply equations for sugar:

Qd =1000-1000p        where Q d is quantity demanded and Qs is quantity supplied. Qs=800+ 1000p

Where P is the price of sugar per pound and Q is thousands of pounds of sugar.

a.  What are the equilibrium price and quantity for sugar? (1point)

b.  Suppose that the government wishes to subsidize sugar production by placing a floor on sugar prices of $0.20 per pound. What would be the relationship between the quantity supplied and quantity demand for sugar?

C. Identify market problem specifically at prices 0.2 per pound and what will be scientific recommendation you suggest to solve the identified market problem?


1
Expert's answer
2021-04-23T12:38:41-0400

(a)

The price and quantity is decided at an equilibrium point, and at Equilibrium Qd  = Qs

"1000-1000p=800+1000p"

"1000-800=1000p+1000p"

"200=2000p"

"P=0.1"

Price of sugar per pound is $0.10

"=1000-1000(0.1)"

"=1000-100"

"Q=900"

Quantity is 900 thousand pounds of sugar.


(b)

Quantity demanded for sugar when priced at $0.20 per pound

"Qd=1000-1000p"

"=1000-1000(0.2)"

"=1000-200"

"Qd=800"

Quantity supplied for sugar when priced at $0.20 per pound

"Qs=800+1000p"

"=800+1000(0.2)"

"=800+200"

"Qs=1000"


At $0.20 per pound of sugar quantity supplied for sugar is more than its quantity demanded.


(c)

The problem with setting the price of sugar at $0.20 is that at this price there will be an excess supply of sugar. As producer will not be able to sell all the quantity of sugar he has produced.

This will force the producer to reduce their prices of sugar in order to make the product more attractive and affordable for the buyers. In response to the reduction in the price, the consumers will increase their quantity demanded as their demand goes up with a fall in the prices of a commodity and producers will not produce sugar as much as they were producing before. The market will eventually become balanced in some time and will reach the equilibrium price and quantity level.